What is the pH when 20 ml 0.1 M HCl is added to 100 ml 0.1M NaCN?


1
Expert's answer
2021-05-17T04:19:57-0400

"HCl+NaCN \Rightarrow HCN+ NaCl"


Now initially the moles of each:

"n_{HCl}=0.02\times0.1=0.002 moles"

"n_{NaCN}=0.1\times0.1=0.01moles"


from above we can see that HCl is a limiting reagent,so it consumes completely.

Hence after the reaction

"n_{HCl}=0"

"n_{NaCN}=" "0.01-0.002=0.008moles"


SO concentration of "NaCN=\dfrac{n_{NaCN}}{Total volume}=\dfrac{0.008}{0.02+0.1}=0.067M"


As "NaCN" is a base so it will gave us "P_{_{OH}}"


Hence "P_{OH}=-log[CN^-]"


"P_{OH}=-log[0.067]=1.17M"


So "P_{H}=14-1.17=12.83M"